CAN U PLZ INTEGRATE 1/COS(x-a)cos(x-b)


 


 


and while registering on askiitians.com i am writing correct emailid robin.kango@yahoo.com but why it is returning a message thet this id is not valid plzzzzzzzzzzz tell me on my id bcoz i want to register on askiitians.com


 

Grade:

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear robim,

I= ∫ { 1 / [ cos ( x - a ). cos ( x - b ) ] } dx

I = csc(a-b) • ∫ { sin(a-b) / [ cos(x-a). cos(x-b) ] } dx


= csc(a-b) • ∫ { sin [ ( x-b) - (x-a) ] / [ cos(x-a).cos(x-b) ] } dx


= csc(a-b) • ∫ { [ sin(x-b).cos(x-a) - cos(x-b).sin(x-a) ] / [ cos(x-a).cos(x-b) ] } dx

..= csc(a-b) • ∫ [ tan(x-b) - tan(x-a) ] dx

..= csc(a-b) • { ln | sec(x-b) | - ln | sec(x-a) | } + C

..= csc(a-b) • ln | sec(x-b) / sec(x-a) | + C

..= csc(a-b) • ln | cos(x-a) / cos(x-b) | + C

 

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

 

Sagar Singh

B.Tech IIT Delhi

vikas askiitian expert
13 years ago

multiply and divide expression by sin(b-a)

 integration of sin(b-a)/sin(b-a)cos(x-a).cos(x-b)

 sin(b-a)=sin{(x-a)-(x-b)}=sinx-a .cosx-b  - cosx-a .sinx-b

  now expression becomes {sinx-a .cosx-b - sinx-b.cosx-a}/sin(b-a).cosx-a .cosx-b

   =1/sin(b-a)integral {tan(x-a) - tan(x-b)}

   =1/sin(b-a). {logsec(x-a)-logsec(x-b)} +c

  =1/sin(b-a). {logsec(x-a)/sec(x-b)} +c